1. if lim x 9 - a 9 /x-a =lim (4+x),find all possible values of a. x->a x->5

1.   if lim x9-a9/x-a =lim (4+x),find all possible values of a.


        x->a                 x->5



Grade:Upto college level

2 Answers

vikas askiitian expert
13 years ago

x9-a9= (x3)3-(a3)3  =(x3-a3)(x6+a6+(ax)3)

                             =(x-a)(x2+a2+ax)(x6+a6+(ax)3)

after putting this ,expression becomes

  lim x->a    (x2+a2+ax)(x6+a6+(ax)3) =  lim x->5  (4+x)

after taking limit

                 (3a2)(3a6)=9

                    a8=1

 a=-1 or 1 a ns

 

Souradeep Majumder
13 years ago

POSSIBLE VALUE OF A IS +1 OR -1.  BECAUSE FOR FIRST LIMIT THATS GONNA BE 9a8 AND FOR SECOND LIMIT ITS GONNA BE 9.  HENCE

 

9a8=9

 

a=+1 AND -1

 

APPROVE MY ANSWER........